Q3. 120 pts) There is some indication that consumption of moderate amounts of wine can help prevent heart attacks. Below is a i need the answers to b c d e and f
Answer #1

Least square formula :- Y = Bo + B1X

B1 =   media%2F060%2F060b989b-a994-4281-adee-a2

Bo = mean of y - B1 * mean of x

ŷ = -14.10801X + 208.28066

b) slope (B1) = -14.10801

c) interpretation of slope :-  slope (B1) = -14.10801 means y decrease -14.10801( per 100000) by increase of one unit x. means heart disease death rate decrease -14.10801 by increase of one unit acchohol consumption( per person).

d) intercept (Bo) = 208.28066

e) interpretation of intercept :- intercept (Bo) = 208.28066 means heart disease death rate is 208.28066 per 100000 when alchohol consumption is minimum per person.

f) equation of least square line :- Y = Bo + B1X

i) spain, wine consumption 6.5 liters per person so, x is 6.5

therefore, Y = 208.28066 +( -14.10801* 6.5)

= 116.578 is predict death rate from heart.

ii) actual death rate from heart disease is 86. so, residual for spain is (116.578 - 86) = 83